OpenStudy (one098):

What is the f(x)/g(x) when f(x) = 6x3 - 19x2 + 16x - 4 and g(x) = x - 2?

OpenStudy (one098):

@Jhannybean

OpenStudy (anonymous):

\[ \frac{f(x)}{g(x)} = \frac{6x^3-19x^2+16x-4}{x-2} \]

OpenStudy (anonymous):

You may want to check if \(x-2\) divides the top.

OpenStudy (anonymous):

\[ f(2) = 6(2)^3-19(2)^2+16(2)-4 = 6(8)-19(4)+16(2)-4 \]

OpenStudy (one098):

So sorry for bothering you wio.

OpenStudy (anonymous):

The reason I do this is if \(f(2) = 0\), then that means \(x-2\) can be factored out.

OpenStudy (anonymous):

Do you know how to divide polynomials?

OpenStudy (one098):

Yes, I think so.

OpenStudy (anonymous):

Okay, first of all, is \(f(2) = 0\)?

OpenStudy (one098):

I dont know, sorry.

OpenStudy (anonymous):

Well, can you try?

OpenStudy (anonymous):

We need to keep simplifying this:\[ f(2) = 6(2)^3-19(2)^2+16(2)-4 = 6(8)-19(4)+16(2)-4 \]

OpenStudy (one098):

Okay. How is it that all these terms dont have the same x value?

OpenStudy (anonymous):

No, they all had \(x=2\), but then I squared/cubed some of them

OpenStudy (anonymous):

They need to be multiplied, so I kept parethesis

OpenStudy (one098):

Oh okay.

OpenStudy (anonymous):

So can you do the operations and find out the final value?

OpenStudy (one098):

Okay, I'll try.

OpenStudy (one098):

I got: \[f(2)=48-76+32-4=48-76+32-4\]

OpenStudy (anonymous):

Okay, now do subtraction and addition to get final result.

OpenStudy (one098):

\[f(2)=-28+28=-28+28\]

OpenStudy (anonymous):

So \(f(2) =0\).

OpenStudy (anonymous):

This means you can simplify if you divide out the \(x-2\). Can you do that?

OpenStudy (one098):

Okay.

OpenStudy (anonymous):

Hmm, what did you get?

OpenStudy (inowalst):

^